OpenStudy (anonymous):

Solve. –3(5 – 15n) = 15(3n – 2)

terenzreignz (terenzreignz):

Distribute.

OpenStudy (anonymous):

danint ur fast

terenzreignz (terenzreignz):

Agreed ;) But what say you, @DUDE..IM..A..DUCK

OpenStudy (mathmale):

@DUDE..IM..A..DUCK : I'm entirely serious. Show your OWN work and do not wait for others to do it for you.

terenzreignz (terenzreignz):

@mathmale In fairness to dud..im..a..duck, in the previous question, I did ask him how he got -4 and he showed legit steps into it...

terenzreignz (terenzreignz):

dude*

OpenStudy (mathmale):

Thank you. Let's see how he does on this particular problem.

OpenStudy (mathmale):

@terenzreignz : I invite you to provide an example of the "distributive property."

terenzreignz (terenzreignz):

Roger. A demonstration of the "distributive property" When a term is multiplied to a sum of terms, such as in the following: \[\Large a ( b + c -d )\] Then the multiplied term is 'distributed', so to speak, to all the terms inside, by multiplying to each term inside the grouping symbol. Like so: \[\Large a(b+c -d) = ab+ac -ad\] Note that your usual rules when it comes to multiplying signs still apply. It is important not to be confused when distributing a negative term. Suppse we have: \[\Large (-7)(2x -9y +3z)\] Then we get \[\Large (-7)(2x) -(-7)(9y) + (-7)(3z)\]\[\Large = -14x + 63y - 21z\] And that's how it's done ^_^

OpenStudy (mathmale):

Really nice work...thank you very much. @DUDE..IM..A..DUCK : Your turn. Please distribute (through multiplication): –3(5 – 15n)

OpenStudy (mathmale):

If you don't "get it," please ask @terenzreignz for guidance.